he nominal value of treasury bill is equal to 10 000 PLN. The price of 90-day...

  1. he nominal value of treasury bill is equal to 10 000 PLN. The price of 90-day treasury bill, which the time to maturity is 60 days, is equal to 9 938 PLN. The interest rate is equal to 3.5% p.a. Is it worth buying this treasury bill? If so, what will be the result on this operation?
Answer #1

Answer to above question:

Alternative 1:

Cost of Treasury bills After 30 Days(Present value of treasury bills)= 10000 -(10000*3.5%*30/365)=9971.24 or 9971 PLN (approx)

Cost at which treasury bills is listed= 9938 PLN

Gain Due to Buying = 9971-9938= 33 PLN

Alternative 2:

Amount which is received after 60 days= 10000+(10000*3.5%*90/365)=10086.30

Cost of Treasury bills=9938

Gain due to buying the treasury bill= 10086.3-9938=148.30